# The Primordial Black Hole Abundance: The Broader, the Better

### v1

@@ 4 The relevance of Broadness: the non-Gaussian case
$\zeta(r)=F[\zeta_{\rm g}(r)].$ (4.1)
In such a case the compaction function is still given by Eq. (2), where
$C_{\zeta}(r)=F_{1}(\zeta_{\rm g})C_{\rm g}(r),\,\,\,\,C_{\rm g}(r)=-\frac{4}{3}r\zeta^{\prime}_{\rm g}(r),$ (4.2)
@@
$C_{\zeta}^{\prime}(r_{m})=F_{1}(\zeta_{\rm g})C^{\prime}_{{\rm g}}(r_{m})+C_{{\rm g}}(r_{m})\zeta_{\rm g}^{\prime}(r_{m})F_{2}(\zeta_{\rm g})=0,$ (4.3)
@@
$x_{0}=\zeta_{\rm g},\quad x_{1}=r\zeta^{\prime}_{\rm g},\quad x_{2}=r^{2}\zeta^{\prime\prime}_{\rm g},\quad x_{3}=r^{3}\zeta^{\prime\prime\prime}_{\rm g},$ (4.4)
@@
$x_{2}=-x_{1}\left(1+x_{1}\frac{F_{2}(x_{0})}{F_{1}(x_{0})}\right).$ (4.5)
@@
$P(x_{0},x_{1},x_{2},x_{3})=\frac{1}{(2\pi)^{2}\sqrt{{\rm det}\,\Sigma}}\,{\rm exp}\left(-\vec{V}^{T}\Sigma^{-1}\vec{V}/2\right),$ (4.6)
@@
$\Sigma$ $=$ $\left(\begin{array}[]{cccc}\sigma^{2}_{0}&\gamma_{01}\sigma_{1}\sigma_{0}&\gamma_{02}\sigma_{2}\sigma_{0}&\gamma_{03}\sigma_{3}\sigma_{0}\\ \gamma_{01}\sigma_{1}\sigma_{0}&\sigma_{1}^{2}&\gamma_{12}\sigma_{2}\sigma_{1}&\gamma_{13}\sigma_{1}\sigma_{3}\\ \gamma_{02}\sigma_{2}\sigma_{0}&\gamma_{12}\sigma_{2}\sigma_{1}&\sigma^{2}_{2}&\gamma_{23}\sigma_{2}\sigma_{3}\\ \gamma_{03}\sigma_{3}\sigma_{0}&\gamma_{13}\sigma_{1}\sigma_{3}&\gamma_{23}\sigma_{2}\sigma_{3}&\sigma_{3}^{2}\end{array}\right)$ (4.11)
@@
$\sigma_{i}^{2}=\langle x_{i}^{2}\rangle,\quad\gamma_{ij}=\frac{\langle x_{i}x_{j}\rangle}{\langle x_{i}^{2}\rangle^{1/2}\;\langle x_{j}^{2}\rangle^{1/2}}.$ (4.12)
@@
$r^{2}C_{\zeta}^{\prime\prime}$ $=$ $-\frac{4}{3}\left[F_{1}(x_{0})(2x_{2}+x_{3})+2x_{1}^{2}F_{2}(x_{0})+3x_{1}x_{2}F_{2}(x_{0})+x_{1}^{3}F_{3}(x_{0})\right].$ (4.14)
@@
$\beta(q)=\int_{D}\mathcal{K}\left(C-C_{c}(q)\right)^{\gamma}\;p(x_{0},C_{\zeta},x_{2},q)\;\delta\left(F_{1}(x_{0})(x_{1}+x_{2})+x_{1}^{2}F_{2}(x_{0})\right),$ (4.15)
@@
$D=\left\{x_{2}\in\mathbb{R},C_{\zeta,c}(q)<C_{\zeta}<\frac{4}{3}\right\},$ (4.16)
@@
$C_{\zeta,c}(q)\simeq\frac{4}{3}\left(1-\sqrt{\frac{2-3C_{c}(q)}{2}}\right).$ (4.17)
We have reintroduced the scaling-law factor for critical collapse $\mathcal{K}(C-C_{c}(q))^{\gamma}$ which accounts for the mass of the PBHs at formation written in units of the horizon mass at the time of horizon re-entry, with $\mathcal{K}\simeq 3.3$ for a log-normal power spectrum and $\gamma\simeq 0.36$ [14, 15, 16] (see also Ref. [17]). By using the conservation of probabilities we can finally write
$p\left[\zeta_{\rm g},C_{\zeta},x_{2},q\right]=P\left[x_{0},x_{1},x_{2},x_{3}\right]|{\rm Det}\,J|,$ (4.18)
@@
${\rm Det}\,J=\frac{3}{4}\left(\frac{4x_{1}+2F_{1}(x_{0})x_{1}^{2}}{1+F_{1}(x_{0})x_{1}}\right),$ (4.19)
@@
$x_{3}=\frac{-4q(1+\frac{1}{2}x_{1}F_{1}(x_{0}))x_{1}}{1+x_{1}F_{1}(x_{0})}-2x_{2}-2x_{1}^{2}\frac{F_{2}(x_{0})}{F_{1}(x_{0})}-3x_{1}x_{2}\frac{F_{2}(x_{0})}{F_{1}(x_{0})}-x_{1}^{3}\frac{F_{3}(x_{0})}{F_{1}(x_{0})}.$ (4.20)
@@
$P(x_{0},x_{1},x_{2},x_{3})=\frac{1}{4\pi^{2}\sqrt{\det{\Sigma}}}\exp{\left(-\frac{(\sigma_{0}\sigma_{1}\sigma_{2}\sigma_{3})^{2}}{2\det{\Sigma}}\sum_{i,j=0}^{3}\frac{\kappa_{ij}x_{i}x_{j}}{\sigma_{i}\sigma_{j}}\right)},$ (4.21)
@@
$A(\zeta_{\rm g},C_{\rm g})=\frac{9216\kappa_{33}C_{\rm g}^{2}F_{1}^{4}\left(8-3C_{\rm g}F_{1}\right){}^{2}}{\sigma_{3}^{2}\left(4-3C_{\rm g}F_{1}\right){}^{2}},$ (4.23)
$B(\zeta_{\rm g},C_{\rm g})$ $=$ $\frac{192C_{\rm g}F_{1}^{2}\left(3C_{\rm g}F_{1}-8\right)}{\sigma_{0}\sigma_{1}\sigma_{2}\sigma_{3}^{2}\left(3C_{\rm g}F_{1}-4\right)}\left[3\kappa_{33}\sigma_{0}\sigma_{1}\sigma_{2}C_{\rm g}\left\{9C_{\rm g}F_{1}\left(C_{\rm g}F_{3}+4F_{2}\right)-27C_{\rm g}^{2}F_{2}^{2}-32F_{1}^{2}\right\}\right.$ (4.24)
@@
4.1 An illustrative example
We consider the following illustrative example which typically arises in models in which the curvature perturbation is generated during a period of ultra-slow-roll [10, 18, 19, 20] [^5] For $\zeta_{\rm g}>\mu_{\star}$, Eq. (4.26) does not capture the possibility of PBHs formed by bubbles of trapped vacuum which requires a separate discussion [21, 22].
$\zeta({\bf x})=-\mu_{\star}\ln\left(1-\frac{\zeta_{\rm g}({\bf x})}{\mu_{\star}}\right),$ (4.26)
with $\mu_{\star}$ a model-dependent parameter depending upon the transition between the ultra-slow-roll phase and the subsequent slow-roll phase. In this analysis we take $\mu_{\star}$ as a free parameter. We take the power spectrum of the Gaussian component to be a log-normal power spectrum
${\cal P}_{\rm g}(k)=\frac{A}{\sqrt{2\pi}\Delta}{\rm exp}\left[-\ln^{2}(k/k_{\star})/2\Delta^{2}\right].$ (4.27)
Our results are summarized in Fig. 4 where we have assumed $k_{\star}r_{m}=2$ for the plot [^6] This is not strictly correct as the value of $r_{m}$ should change for each profile, but it is nevertheless in the ballpark of the value we have assumed [8].. The broadness of the power spectrum is controlled by the parameter $\Delta$. We observe that by increasing the value of $\Delta$, enlarging the power spectra, again the PBHs formation probability is dominated by the broadest profiles. We have checked that for very peaked power spectrum, as in the case for $\Delta=1/3$, the abundance is peaked again around the average of $q$.
Figure 4: Mass fraction $\beta$ for the non-Gaussian scenario computed with several values of $\Delta$, where we fix $\mu_{*}=5/2$, $k_{*}r_{m}=2$ and the amplitude of the power spectrum $A=10^{-2}$.
5 Comparison with literature and impact on the physics of PBHs and Pulsar Timing Arrays
In this section we compare the calculation presented above, accounting for the curvature of the compaction function at its peak, with the prescription based on threshold statistics on the compaction function, reported in Refs. [23, 24], where the only explicit dependence on $q$ is encoded in $C_{c}(q)$. There, the formation probability is computed by integrating the joint probability distribution function $P_{\mathrm{g}}$
$\beta=\int_{D}\mathcal{K}\left(C-C_{c}(q)\right)^{\gamma}\textrm{P}_{\rm g}(C_{\rm g},\zeta_{\rm g}){\rm d}C_{\rm g}{\rm d}\zeta_{\rm g}\,,$ (5.1)
where the domain of integration is given by $\mathcal{D}=\left\{C(C_{\rm g},\zeta_{\rm g})>C_{\rm c}(q),~{}C_{\zeta}(C_{\rm g},\zeta_{\rm g})<4/3\right\}$. The Gaussian components are distributed as
$P_{\rm g}\left(C_{\rm g},\zeta_{\rm g}\right)=\frac{1}{2\pi\sigma_{\rm a}\sigma_{\rm c}\sqrt{1-\gamma_{*}^{2}}}\exp\left[-\frac{1}{2\left(1-\gamma_{*}^{2}\right)}\left(\frac{C_{\mathrm{g}}}{\sigma_{\rm a}}-\frac{\gamma_{*}\zeta_{\mathrm{g}}}{\sigma_{\rm c}}\right)^{2}\!-\!\frac{\zeta_{\mathrm{g}}^{2}}{2\sigma_{\rm c}^{2}}\right],$ (5.2)
@@ 5 Comparison with literature and impact on the physics of PBHs and Pulsar Timing Arrays
$\langle C^{2}_{\rm g}\rangle$ $=\sigma_{\rm a}^{2}=\frac{16}{81}\int_{0}^{\infty}\frac{{\rm d}k}{k}\left(kr_{m}\right)^{4}W^{2}\left(k,r_{m}\right)T\left(k,r_{m}\right)P_{\zeta},$ (5.3a)
$\langle C_{\rm g}\zeta_{\rm g}\rangle$ $=\sigma_{\rm b}^{2}=\frac{4}{9}\!\int_{0}^{\infty}\!\!\frac{{\rm d}k}{k}\!\left(kr_{m}\right)^{2}\!W\!\!\left(k,r_{m}\right)\!W_{s}\!\left(k,r_{m}\right)T\left(k,r_{m}\right)\!P_{\zeta},$ (5.3b)
$\langle\zeta^{2}_{\rm g}\rangle$ $=\sigma_{\rm c}^{2}=\int_{0}^{\infty}\frac{{\rm d}k}{k}W_{s}^{2}\left(k,r_{m}\right)T\left(k,r_{m}\right)P_{\zeta},$ (5.3c)
and $\gamma_{*}=\sigma_{\rm b}^{2}/\sigma_{\rm a}\sigma_{\rm c}.$ We have defined $W\left(k,r_{m}\right)$ and $W_{s}\left(k,r_{m}\right)$ as the top-hat window function and the spherical-shell window function [25]. To compare this prescription with the one presented in this paper, we consider two cases: $\beta_{0}$, in which we do not adopt any transfer function ($T=1$) since everything is determined on superhorizon scales, and $\beta_{T}$, in which we consider the radiation transfer function assuming a perfect radiation fluid, as adopted in Ref. [23].
In Fig. 5, we show a comparison between the two prescriptions using the typical non-Gaussian relation in the ultra-slow-roll scenario (see Eq. (4.26)) with a log-normal power spectrum (see Eq. (4.27)) with several benchmark values for $\mu_{\star}$. We fix $\Delta=1$ in the plots, but we have found analogous results also varying this parameter. As we can understand from Fig. 5, evaluating the quantities on superhorizon scales, i.e. the ratio $\beta/\beta_{0}$, the old prescription marginally overestimates the abundance of PBHs. This discrepancy arises because, unlike the prescription used in the literature, where an average profile is employed, the effective threshold is slightly larger than the averaged case, as evident from Eq. (3.23). Nevertheless an equivalent amount of PBHs requires a slightly larger amplitude of the curvature perturbation power spectrum.
@@
This has important implications for the phenomenology related to PBHs respect to the case discussed in this paper. Indeed in the standard formation scenario PBH formation occurs as large curvature perturbations re-enter the Hubble horizon after inflation and eventually collapse under the effect of gravity. When such scalar perturbations cross the horizon they produce tensor modes as a second-order effect, which appear to us today as a signal of stochastic gravitational waves background (SGWB) (for a recent review, see Ref. [27]). Recently in Ref.[26], where the old prescription was used, it was shown that large negative non-Gaussianity are necessary in order to achieve high enough amplitude, without overproducing PBHs, in order to relax the tension between the PTA recent dataset (the most constrained dataset is the one released by NANOGrav [28]) and the PBH explanation.
We demonstrate that even when correctly accounting for the impact of the curvature of the compaction function and calculating all the relevant quantities on superhorizon scales, thereby avoiding all concerns regarding non-linearities in the radiation transfer function and the determination of the true physical horizon, the tension between the PTA dataset and the PBH hypothesis is even worse than what claimed in Ref. [26].
Figure 5: Ratio between mass fraction $\beta$ for the non-Gaussian case between the prescriptions presented in this paper and the prescription presented in Ref. [23]. We fix the shape parameter $q=0.5$ (as a consequence also the threshold using Eq. (2.11)) and the shape of power spectrum $\Delta=1$ while we vary the amplitude.
6 Conclusions and some further final considerations

### v2

@@ IV The relevance of Broadness: the non-Gaussian case
$\zeta(r)=F[\zeta_{\rm g}(r)].$ (41)
In such a case the compaction function is still given by Eq. (II), where
$C_{\zeta}(r)=F_{1}(\zeta_{\rm g})C_{\rm g}(r),\,\,\,\,C_{\rm g}(r)=-\frac{4}{3}r\zeta^{\prime}_{\rm g}(r),$ (42)
@@
$C_{\zeta}^{\prime}(r_{m})=F_{1}(\zeta_{\rm g})C^{\prime}_{{\rm g}}(r_{m})+C_{{\rm g}}(r_{m})\zeta_{\rm g}^{\prime}(r_{m})F_{2}(\zeta_{\rm g})=0,$ (43)
@@
$x_{0}=\zeta_{\rm g},\quad x_{1}=r\zeta^{\prime}_{\rm g},\quad x_{2}=r^{2}\zeta^{\prime\prime}_{\rm g},\quad x_{3}=r^{3}\zeta^{\prime\prime\prime}_{\rm g},$ (44)
@@
$x_{2}=-x_{1}\left(1+x_{1}\frac{F_{2}(x_{0})}{F_{1}(x_{0})}\right).$ (45)
@@
$P(x_{0},x_{1},x_{2},x_{3})=\frac{1}{(2\pi)^{2}\sqrt{{\rm det}\,\Sigma}}\,{\rm exp}\left(-\vec{V}^{T}\Sigma^{-1}\vec{V}/2\right),$ (46)
@@
$\Sigma$ $=$ $\left(\begin{array}[]{cccc}\sigma^{2}_{0}&\gamma_{01}\sigma_{1}\sigma_{0}&\gamma_{02}\sigma_{2}\sigma_{0}&\gamma_{03}\sigma_{3}\sigma_{0}\\ \gamma_{01}\sigma_{1}\sigma_{0}&\sigma_{1}^{2}&\gamma_{12}\sigma_{2}\sigma_{1}&\gamma_{13}\sigma_{1}\sigma_{3}\\ \gamma_{02}\sigma_{2}\sigma_{0}&\gamma_{12}\sigma_{2}\sigma_{1}&\sigma^{2}_{2}&\gamma_{23}\sigma_{2}\sigma_{3}\\ \gamma_{03}\sigma_{3}\sigma_{0}&\gamma_{13}\sigma_{1}\sigma_{3}&\gamma_{23}\sigma_{2}\sigma_{3}&\sigma_{3}^{2}\end{array}\right)$ (51)
@@
$\sigma_{i}^{2}=\langle x_{i}^{2}\rangle,\quad\gamma_{ij}=\frac{\langle x_{i}x_{j}\rangle}{\langle x_{i}^{2}\rangle^{1/2}\;\langle x_{j}^{2}\rangle^{1/2}}.$ (52)
@@
$r^{2}C_{\zeta}^{\prime\prime}$ $=$ $-\frac{4}{3}\left[F_{1}(x_{0})(2x_{2}+x_{3})+2x_{1}^{2}F_{2}(x_{0})+3x_{1}x_{2}F_{2}(x_{0})+x_{1}^{3}F_{3}(x_{0})\right].$ (54)
@@
$\beta(q)=\int_{D}\mathcal{K}\left(C-C_{c}(q)\right)^{\gamma}\;p(x_{0},C_{\zeta},x_{2},q)\;\delta\left(F_{1}(x_{0})(x_{1}+x_{2})+x_{1}^{2}F_{2}(x_{0})\right),$ (55)
@@
$D=\left\{x_{2}\in\mathbb{R},C_{\zeta,c}(q)<C_{\zeta}<\frac{4}{3}\right\},$ (56)
@@
$C_{\zeta,c}(q)\simeq\frac{4}{3}\left(1-\sqrt{\frac{2-3C_{c}(q)}{2}}\right).$ (57)
We have reintroduced the scaling-law factor for critical collapse $\mathcal{K}(C-C_{c}(q))^{\gamma}$ which accounts for the mass of the PBHs at formation written in units of the horizon mass at the time of horizon re-entry, with $\mathcal{K}\simeq 3.3$ for a log-normal power spectrum and $\gamma\simeq 0.36$ Choptuik_gamma ; crit_1 ; crit_2 ; Ianniccari:2024ltb (see also Ref. Sasaki_review ). By using the conservation of probabilities we can finally write
$p\left[\zeta_{\rm g},C_{\zeta},x_{2},q\right]=P\left[x_{0},x_{1},x_{2},x_{3}\right]|{\rm Det}\,J|,$ (58)
@@
${\rm Det}\,J=\frac{3}{4}\left(\frac{4x_{1}+2F_{1}(x_{0})x_{1}^{2}}{1+F_{1}(x_{0})x_{1}}\right),$ (59)
@@
$x_{3}=\frac{-4q(1+\frac{1}{2}x_{1}F_{1}(x_{0}))x_{1}}{1+x_{1}F_{1}(x_{0})}-2x_{2}-2x_{1}^{2}\frac{F_{2}(x_{0})}{F_{1}(x_{0})}-3x_{1}x_{2}\frac{F_{2}(x_{0})}{F_{1}(x_{0})}-x_{1}^{3}\frac{F_{3}(x_{0})}{F_{1}(x_{0})}.$ (60)
@@
$P(x_{0},x_{1},x_{2},x_{3})=\frac{1}{4\pi^{2}\sqrt{\det{\Sigma}}}\exp{\left(-\frac{(\sigma_{0}\sigma_{1}\sigma_{2}\sigma_{3})^{2}}{2\det{\Sigma}}\sum_{i,j=0}^{3}\frac{\kappa_{ij}x_{i}x_{j}}{\sigma_{i}\sigma_{j}}\right)},$ (61)
@@
$A(\zeta_{\rm g},C_{\rm g})=\frac{9216\kappa_{33}C_{\rm g}^{2}F_{1}^{4}\left(8-3C_{\rm g}F_{1}\right){}^{2}}{\sigma_{3}^{2}\left(4-3C_{\rm g}F_{1}\right){}^{2}},$ (63)
$B(\zeta_{\rm g},C_{\rm g})$ $=$ $\frac{192C_{\rm g}F_{1}^{2}\left(3C_{\rm g}F_{1}-8\right)}{\sigma_{0}\sigma_{1}\sigma_{2}\sigma_{3}^{2}\left(3C_{\rm g}F_{1}-4\right)}\left[3\kappa_{33}\sigma_{0}\sigma_{1}\sigma_{2}C_{\rm g}\left\{9C_{\rm g}F_{1}\left(C_{\rm g}F_{3}+4F_{2}\right)-27C_{\rm g}^{2}F_{2}^{2}-32F_{1}^{2}\right\}\right.$ (64)
@@
IV.1 An illustrative example
We consider the following illustrative example which typically arises in models in which the curvature perturbation is generated during a period of ultra-slow-roll cai ; b1 ; b2 ; Tomberg:2023kli  [^5] For $\zeta_{\rm g}>\mu_{\star}$, Eq. (66) does not capture the possibility of PBHs formed by bubbles of trapped vacuum which requires a separate discussion Escriva:2023uko ; Uehara:2024yyp .
$\zeta({\bf x})=-\mu_{\star}\ln\left(1-\frac{\zeta_{\rm g}({\bf x})}{\mu_{\star}}\right),$ (66)
with $\mu_{\star}$ a model-dependent parameter depending upon the transition between the ultra-slow-roll phase and the subsequent slow-roll phase. To focus only on the impact of primordial non gaussianity, in this analysis we take $\mu_{\star}$ as a free parameter. We take the power spectrum of the Gaussian component to be a log-normal power spectrum
${\cal P}_{\rm g}(k)=\frac{A}{\sqrt{2\pi}\Delta}{\rm exp}\left[-\ln^{2}(k/k_{\star})/2\Delta^{2}\right].$ (67)
Our results are summarized in Fig. 4 where, changing the shape of the power spectrum, we computed $k_{\star}r_{m}$ following ref musconl  [^6] Here we stress that the value of $r_{m}$ can slighlty change in presence of large primordial non-guassianities Kehagias:2019eil .. The broadness of the power spectrum is controlled by the parameter $\Delta$. We observe that by increasing the value of $\Delta$, enlarging the power spectra, again the PBHs formation probability is dominated by the broadest profiles. We have checked that for very peaked power spectrum, as in the case for $\Delta=1/3$, the abundance is peaked again around the average of $q$.
Figure 4: Mass fraction $\beta$ for the non-Gaussian scenario computed with several values of $\Delta$, where we fix $\mu_{*}=5/2$ and the amplitude of the power spectrum $A=10^{-2}$.
V Comparison with literature and impact on the physics of PBHs and Pulsar Timing Arrays
In this section we compare the calculation presented above, accounting for the curvature of the compaction function at its peak, with the prescription based on threshold statistics on the compaction function, reported in Refs. F ; Gow , where the only explicit dependence on $q$ is encoded in $C_{c}(q)$. There, the formation probability is computed by integrating the joint probability distribution function $P_{\mathrm{g}}$
$\beta=\int_{D}\mathcal{K}\left(C-C_{c}(q)\right)^{\gamma}\textrm{P}_{\rm g}(C_{\rm g},\zeta_{\rm g}){\rm d}C_{\rm g}{\rm d}\zeta_{\rm g}\,,$ (68)
where the domain of integration is given by $D=\left\{C(C_{\rm g},\zeta_{\rm g})>C_{\rm c}(q),\leavevmode\nobreak\ C_{\zeta}(C_{\rm g},\zeta_{\rm g})<4/3\right\}$. The Gaussian components are distributed as
$P_{\rm g}\left(C_{\rm g},\zeta_{\rm g}\right)=\frac{1}{2\pi\sigma_{\rm a}\sigma_{\rm c}\sqrt{1-\gamma_{*}^{2}}}\exp\left[-\frac{1}{2\left(1-\gamma_{*}^{2}\right)}\left(\frac{C_{\mathrm{g}}}{\sigma_{\rm a}}-\frac{\gamma_{*}\zeta_{\mathrm{g}}}{\sigma_{\rm c}}\right)^{2}\!-\!\frac{\zeta_{\mathrm{g}}^{2}}{2\sigma_{\rm c}^{2}}\right],$ (69)
@@ V Comparison with literature and impact on the physics of PBHs and Pulsar Timing Arrays
$\langle C^{2}_{\rm g}\rangle$ $=\sigma_{\rm a}^{2}=\frac{16}{81}\int_{0}^{\infty}\frac{{\rm d}k}{k}\left(kr_{m}\right)^{4}W^{2}\left(k,r_{m}\right)T^{2}\left(k,r_{m}\right)P_{\zeta},$ (70a)
$\langle C_{\rm g}\zeta_{\rm g}\rangle$ $=\sigma_{\rm b}^{2}=\frac{4}{9}\!\int_{0}^{\infty}\!\!\frac{{\rm d}k}{k}\!\left(kr_{m}\right)^{2}\!W\!\!\left(k,r_{m}\right)\!W_{s}\!\left(k,r_{m}\right)T^{2}\left(k,r_{m}\right)\!P_{\zeta},$ (70b)
$\langle\zeta^{2}_{\rm g}\rangle$ $=\sigma_{\rm c}^{2}=\int_{0}^{\infty}\frac{{\rm d}k}{k}W_{s}^{2}\left(k,r_{m}\right)T^{2}\left(k,r_{m}\right)P_{\zeta},$ (70c)
and $\gamma_{*}=\sigma_{\rm b}^{2}/\sigma_{\rm a}\sigma_{\rm c}.$ We have defined $W\left(k,r_{m}\right)$ and $W_{s}\left(k,r_{m}\right)$ as the top-hat window function and the spherical-shell window function Young . To compare this prescription with the one presented in this paper, we consider two cases: $\beta_{0}$, in which we do not adopt any transfer function ($T=1$) since everything is determined on superhorizon scales, and $\beta_{T}$, in which we consider the radiation transfer function assuming a perfect radiation fluid, as adopted in Ref. F .
In Fig. 5, we show a comparison between the two prescriptions using the typical non-Gaussian relation in the ultra-slow-roll scenario (see Eq. (66)) with a log-normal power spectrum (see Eq. (67)) with several benchmark values for $\mu_{\star}$. We fix $\Delta=1$ in the plots, but we have found analogous results also varying this parameter. As we can understand from Fig. 5, evaluating the quantities on superhorizon scales, i.e. the ratio $\beta/\beta_{0}$, there is a marginal discrepancy between the two prescription. This discrepancy arises because, unlike the prescription used in the literature, where an average profile is employed, the effective threshold is slightly different than the averaged case, as evident from Eq. (39). Nevertheless an equivalent amount of PBHs requires a marginal change in the amplitude of the curvature perturbation power spectrum.
@@
This has important implications for the phenomenology related to PBHs respect to the case discussed in this paper. Indeed in the standard formation scenario PBH formation occurs as large curvature perturbations re-enter the Hubble horizon after inflation and eventually collapse under the effect of gravity. When such scalar perturbations cross the horizon they produce tensor modes as a second-order effect, which appear to us today as a signal of stochastic gravitational waves background (SGWB) (for a recent review, see Ref. Domenech:2021ztg ). Recently in Ref.IovinoPTA , where the old prescription was used, it was shown that large negative non-Gaussianity are necessary in order to achieve high enough amplitude, without overproducing PBHs, in order to relax the tension between the PTA recent dataset (the most constrained dataset is the one released by NANOGrav NANOGrav ) and the PBH explanation.
We demonstrate that even when correctly accounting for the impact of the curvature of the compaction function and calculating all the relevant quantities on superhorizon scales, thereby avoiding all concerns regarding non-linearities in the radiation transfer function and the determination of the true physical horizon, the tension between the PTA dataset and the PBH hypothesis is even worse than what claimed in Ref. IovinoPTA .
Figure 5: Ratio between mass fraction $\beta$ for the non-Gaussian case between the prescriptions presented in this paper and the prescription presented in Ref. F . We fix the shape parameter $q=0.5$ (as a consequence also the threshold using Eq. (11)) and the shape of power spectrum $\Delta=0.5$ while we vary the amplitude.
We conclude this section, making a general comparison with another statistical approach for computing the PBH abundance. When the abundance is exponentially sensitive on the threshold, as well as in the case of peak theory Bardeen:1985tr ; GreenPeak ; Yoo:2018kvb ; Yoo:2019pma ; Yoo:2020dkz ; Young:2020xmk ; Kitajima:2021fpq ; Taoso:2021uvl ; Riccardi:2021rlf ; Young (see for example section 3 of ref. Yoo:2019pma for a general expression of the PBH fraction in the context of peak theory or simply Eq. (32) and Eq. (33) of ref. Young ) we expect that our results can be generally extended to these other approaches. However there are still discrepancies between these two approaches, which are already present at Gaussian level. Indeed the approach based on peak theory requires slightly smaller values of the amplitude in order to get the same abundance of PBHs Young1 ; DeLuca1 , thus making the claim on the tension with the PTA dataset, even stronger. We leave a deeper analysis for the discrepancies between threshold statistics and peak theory in presence of primordial non-gaussianities for a future work.
VI Conclusions and some further final considerations